OpenStudy (anonymous):

the domain and range of y=square root of x^2 -4

OpenStudy (anonymous):

Domain = (x:x greater or equal to -4, x is any real number) Range= (y:y is any real number)

OpenStudy (anonymous):

for domain \[x^{2}-4\ge 0, x ^{2} \ge4 \] \[\left| x \right|\ge 2, x \le-2 and x \ge 2\] \[Domain=\left\{ x:x \le -2 or x \ge 2 \right\}\] \[Range=\left\{ y:y \ge 0 \right\}\]